The company with the common equity accounts shown here has decided on a two-for-one stock split. The firm’s 34-cent-per-share cash dividend on the new (postsplit) shares represents an increase of 10 percent over last year’s dividend on the presplit stock.

  Common stock ($1 par value) $   500,000
  Capital surplus 1,558,000
  Retained earnings 3,884,000
  Total owners’ equity $ 5,942,000
a. What is the new par value of the stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
b. What was last year’s dividend per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

a. New par value per share ___

b. Last years dividend per share ___
Answer #1

New par value: New par value Calculating based on ratio of old share to new share

old share par value is 1, after decided on a two for one stock split then new share will be

New share par value= old share par value ×1/2= 1×1/2

New share par value= 0.50 per share

Calculation of last year dividend per share

First find out the current year dividend per share :

Common stock par Total value is 500,000 with par value 1 so that outstanding share is 500,000

Current year dividend = number of outstanding share × Firm share cash dividend × stock split

Current year total dividend = 500,000 × 34% ×2/1

Current year total dividend =340,000

Last year total dividend = current year total dividend × 10% increase in last year (110%)

Last year total dividend = 340000. × 110% =374000

Last year dividend per share= last year total dividend / number of outstanding share= 374000/500000

Last dividend per share= 0.748